OpenStudy (anonymous):

What is the mass, in grams, of a sample of 1.72 × 1023 atoms of potassium (K)? Show your work or explain the steps that you used to determine your answer.

OpenStudy (mertsj):

\[1.72 \times10^{23}atoms K \times\frac{1 mole K}{6.023\times10^{23}atoms K}\times\frac{39.098 g K}{1 mole K}\]